OpenStudy (anonymous):

Convert the following parametric equations into Cartesian equations. x = cos2A y = sin^2 A

OpenStudy (anonymous):

not sure how to proceed with this it might be useful to use the identity cos 2A = 1 - 2 sin^A

OpenStudy (anonymous):

no its cos 2A = 1 - 2 sin^2 A

OpenStudy (anonymous):

so 1 - 2 sin^2 A = 1 - 2y s0 x = 1 - 2y

OpenStudy (anonymous):

Thanks for your help! I've just attempted using 1-2sin^2 A and i've come out with the answer: (1-x)/2 = sin^2A which makes sense as then I was able to replace y = sin^2 A with y = (1-x)/2 <-- the answer. Thanks again :)

OpenStudy (anonymous):

in standard form thats x + 2y = 1 right - good work